OpenStudy (anonymous):

How to find the missing number in this series 4 , 6 , 9 , 1.8 , 3.8 , ...... , 1.36

OpenStudy (anonymous):

It looks like there are sets of 3... First set increases by 2 and then 3... (i.e. 4 + 2 = 6 + 3 = 9) Do the same for the second set... 1.8 + 2 = 3.8 + 3 = 6.8 1.36 starts the third set